OpenStudy (anonymous):

How is lim x-> pi/2 from the right = -infinity?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \frac{ \pi }{ 2 }^+}\frac{ 1 }{ 0 }=-\infty \]

OpenStudy (anonymous):

The original problem is this:\[\lim_{x \rightarrow \frac{ \pi }{ 2 }^+}\frac{ cosx }{ 1-sinx }\]